### v1

# STGFormer: Spatio-Temporal GraphFormer for 3D Human Pose Estimation in Video

Abstract
The current methods of video-based 3D human pose estimation have achieved significant progress; however, they continue to confront the significant challenge of depth ambiguity. To address this limitation, this paper presents the spatio-temporal GraphFormer framework for 3D human pose estimation in video, which integrates body structure graph-based representations with spatio-temporal information. Specifically, we develop a spatio-temporal criss-cross graph (STG) attention mechanism. This approach is designed to learn the long-range dependencies in data across both time and space, integrating graph information directly into the respective attention layers. Furthermore, we introduce the dual-path modulated hop-wise regular GCN (MHR-GCN) module, which utilizes modulation to optimize parameter usage and employs spatio-temporal hop-wise skip connections to acquire higher-order information. Additionally, this module processes temporal and spatial dimensions independently to learn their respective features while avoiding mutual influence. Finally, we demonstrate that our method achieves state-of-the-art performance in 3D human pose estimation on the Human3.6M and MPI-INF-3DHP datasets.
Index Terms:
3D human pose estimation, vision transformer, graph convolution
I Introduction
3D human pose estimation predicts the coordinates of body keypoints in three dimensions from input images or videos. Compared to 2D pose estimation, it provides more comprehensive and accurate spatial information. In the field of computer vision, this technology has garnered widespread attention due to its potential applications in action recognition [1, 2], human-computer interaction [3], and autonomous driving [4, 5]. 3D human pose estimation can primarily be divided into single-stage methods and two-stage methods. Single-stage methods directly estimate the end-to-end 3D human pose from inputs; two-stage methods initially detect the coordinates of keypoints in images and then regress to a 3D human pose based on the detected 2D keypoints. Currently, research shows that the performance metrics of two-stage methods generally surpass those of single-stage methods, mainly due to the significant advancements in 2D pose estimation networks. For instance, ViTPose [6] utilizes a transformer as the backbone to extract features and employs a lightweight decoder for pose estimation. The proposed method has been demonstrated to be effective and flexible in 2D pose estimation. However, the second stage, which involves transitioning from 2D pose to 3D, remains unsatisfactory due to the ill-posed problem. This issue stems from the depth ambiguity inherent in estimating 3D coordinates from 2D, caused by a lack of depth prior information.
To address the aforementioned issues, a series of methods have been proposed that are based on backbones such as CNNs [7, 8, 9, 10], GCNs [11, 12, 13, 14, 15], and transformers [16, 17, 18, 19, 20, 21, 22]. For example, Einfalt et al. [19] perform temporal upsampling within Transformer blocks to process time-sparse 2D pose sequences, thereby reducing the model’s size while still generating time-dense 3D pose estimates. StridedFormer [22] reduces sequence redundancy by substituting the fully connected layers in the feedforward network of the Vanilla Transformer Encoder (VTE) with strided convolutions, thereby progressively decreasing the sequence length to lower computational costs. Although the use of transformers capitalizes on the advantages of scalable model capacity and high parallelism, which show promising prospects in enhancing the performance of pose estimation, several challenges still persist.
First, owing to the absence of structural priors in the transformer, its performance in 3D human pose estimation still trails behind the state-of-the-art methods based on GCNs. These graph-based methods effortlessly represent the human skeleton as a graph, where joints are regarded as nodes, and the bones between joints are considered as edges connecting these nodes. Currently, some improvements that integrate graphs with transformers have emerged [23, 24, 25]. For instance, STCFormer [25] introduces body structure graph information as inductive biases into the transformers. This strategy influences the versatility of the transformers, which are commonly regarded as structures with weak inductive biases, with significant implications for pre-training and the scalability of large models.
Second, spatio-temporal transformers excel in capturing long-range dependencies, whereas GCN architectures integrate information and inter-node relationships. Encoding graph information into the attention layer significantly enhances the model’s understanding of direct connections between nodes. Meanwhile, replacing MLP with GCN in transformer may amplifie the ability to aggregate information throughout the entire graph structure. Integrating these two approaches allows the model to process graph data with greater precision and efficiency, thus optimizing performance from different aspects. However, in the most of GCN architectures, the transformation matrices are shared among all nodes within the multi-layer convolutional filters, impeding the efficiency of information exchange among nodes. To address this issue, mechanisms such as weight unsharing [26] and weight modulation [27] have been introduced in the spatial dimension, yet these solutions have not been effectively extended to the temporal dimension.
Third, higher-order dependencies are often neglected in the processing of graph neural networks. Although some studies address these dependencies in the spatial dimension by extending the first-order neighbors to high-order neighbors [28, 29], they remain unresolved in the temporal dimension. This issue of temporal higher-order dependencies originates from the system’s dynamic characteristics, whereby the current state is influenced not only by the immediately preceding state but also by earlier states.
To overcome the limitations discussed above, this paper presents a framework that integrates spatio-temporal transformer with graph encoding and dual-path modulated hop-wise regular GCN for 3D human pose estimation. This framework not only captures spatio-temporal long-range dependencies but also elucidates the relationships between spatio-temporal nodes and their structures. We summarize the main contributions of this work as follows.
(1) We propose a spatio-temporal GraphFormer for 3D human pose estimation from monocular video using spatio-temporal criss-cross graph attention and dual-path modulated hop-wise regular GCN.
(2) We introduce a spatio-temporal criss-cross graph attention mechanism which encodes temporal and spatial graph structural relationships into attention layers.
(3) We design a dual-path modulated hop-wise regular GCN module and integrate it into our GraphFormer architecture.
(4) We demonstrate extensive experiments conducted on Human3.6M and MPI-INF-3DHP datasets and ablation studies that our proposed model achieves state-of-the-art performance in comparison with strong baselines.
II Related Work
In recent years, graph convolutional networks and transformers have made significant progress in 3D human pose estimation. As previously mentioned, we implement 3D human pose estimation from image sequences or video inputs through a two-stage approach.
II-A GCN-based methods.
The skeletal structure of body can be naturally represented as a graph, hence there has been extensive research in image-based 3D human pose estimation using GCNs, which has yielded significant results [13, 15, 26, 27, 30, 31]. To more effectively address the ill-posed problem of lifting 2D joint keypoints into 3D space, additional temporal information is extracted from sequential frames combined with spatial information related to body structure. For instance, the Global-local Adaptive Graph Convolutional Network (GLA-GCN) [11] models the spatio-temporal structure using a graph representation and traces local joint features through individually connected layers. Zeng et al. [32] developed a hop-aware hierarchical channel-squeezing fusion mechanism to efficiently extract relevant information from neighboring nodes in GCNs. Hu et al. [33] represent the human skeleton as a directed graph to effectively reflect the hierarchical relationships among the nodes. Although graph-based methods effectively utilize the body’s structural information, they encounter a significant limitation in terms of scalability. Adding more network layers does not enhance model performance, and even widely adopted residual connections yield minimal impact.
II-B Transformer-based methods.
In addition to GCN, transformer also used to model spatio-temporal correlations for video-based 3D human pose estimation [17, 18, 19, 20, 21, 22, 34, 35]. Specifically, MixSynthFormer [17] combines synthesized spatial and temporal attention to effectively integrate inter-joint and inter-frame significance. HSTFormer [18] captures multi-level joint spatial-temporal correlations, ranging from local to global, through four transformer encoders and a fusion module. Li et al. [22] utilized a vanilla transformer encoder to model long-range dependencies in 2D pose sequences, while a strided transformer encoder aggregates this long-range information into a single-vector representation, employing a hierarchical global and local approach.
The Transformer has high scalability and is good at capturing long-distance dependencies, but it lacks structural prior information. As a result, several methods have been developed that integrate graph and transformer. The Spatio-Temporal Criss-cross Attention (STC) [25] learns spatio-temporal features and incorporates the body’s graph structure information as a biased input in the proposed framework. Zhai et al. [24] proposed the Hop-wise GraphFormer, which enhances GCN-based pose estimation networks by grouping joints according to k-hop neighborhoods and capturing potential joint correlations across various joint synergies. Zhao et al. [23] designed a novel model, GraFormer, that combines graph convolution and transformers to exploit relationships among graph-structured 2D joints.
Our work also belongs to the realm of 3D pose lifting that employs a fusion of transformer and graph. Unlike other approaches, our model integrates graph information directly into the attention layers, thereby significantly improving adaptability. Furthermore, it incorporates spatio-temporal hop-wise modulated GCNs within the transformer framework, effectively capturing high-order dependencies in both time and space. This approach complements the long-range dependencies typical of transformers, thus substantially enhancing the model’s capacity to represent complex data structures.
III Method
In this section, we first clarify the objectives of our learning task. Then, we present the main components of the proposed spatio-temporal graphformer for 3D human pose estimation, including the spatio-temporal criss-cross graph attention and the dual-path modulated hop-wise regular GCN module.
III-A Preliminary
First, let us clarify the problem that needs to be solved in 3D human estimation based on deep learning. The training set for 3D human pose estimation contains paired data $\mathcal{D}={(\mathbf{x}_{i},\mathbf{y}_{i})},{i=1,\cdots,N}$, including 2D joint positions $\mathbf{X}=(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})^{\mathsf{T}}\in\mathbb{R}^{N\times 2}$ and their associated 3D joint positions $\mathbf{Y}=(\mathbf{y}_{1},\cdots,\mathbf{y}_{N})^{\mathsf{T}}\in\mathbb{R}^{N\times 3}$. The goal of human pose lifting is to minimize the loss function $l(f(\mathbf{x}_{i}),\mathbf{y}_{i})$ of a regression model $f:\mathbf{X}\rightarrow\mathbf{Y}$ by learning the parameters.
Then, we provide some preliminary knowledge regarding human pose estimation using graph learning methods. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}={1,\cdots,N}$ denotes the set of $N$ nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges. In the context of 3D human pose estimation, nodes correspond to human joints, and edges represent the connections between these joints. The adjacency matrix $\mathbf{A}=(a_{ij})$, with dimensions $N\times N$, contains elements $a_{ij}$ that represent the weights between node $i$ and node $j$. These weights may be expressed as binary or real numbers. In our approach, we utilize a binary representation, where a value of 1 indicates a connection $i\sim j$ between joints, while a value of 0 signifies the absence of such a connection.
The feature processing can transform low-dimensional joints into high-dimensional features with a dimension of $F$. Each joint can be represented by an $N\times F$ feature matrix $\mathbf{H}=(\mathbf{h}_{1},\cdots,\mathbf{h}_{N})^{\mathsf{T}}$, where each $\mathbf{h}_{i}$ is an $F$-dimensional feature vector corresponding to node $i$. For the joint inputs across consecutive frames $T$, they can be represented by a $T\times N\times F$ feature matrix $\mathbf{H}\in\mathbb{R}^{T\times N\times F}$. In the vanilla GCN [36], a graph convolutional layer transforms and aggregates the input features according to the equation below:
$\mathbf{H}^{\prime}=\sigma(\mathbf{WH}\tilde{\mathbf{A}})$ (1)
where the $i$-th row of the output feature matrix can be expressed as follows:
$\mathbf{h}_{i}^{\prime}=\sigma(\sum_{j\in\mathcal{N}_{i}}{\mathbf{Wh}_{j}}\tilde{a}_{ij})$ (2)
But the shared feature transformation weight $\mathbf{W}$ in the vanilla GCN restricts its ability to learn diverse relational patterns from different body joint nodes. Liu et al. [26] address this limitation by utilizing a unique weight matrix $\mathbf{W}_{j}\in\mathbb{R}^{F^{\prime}\times F}$ for each node $j$, as represented by the following equation:
$\mathbf{h}_{i}^{\prime}=\sigma(\sum_{j\in\mathcal{N}_{i}}{\mathbf{W}_{j}}\mathbf{h}_{j}\tilde{a}_{ij})$ (3)
To reduce the model size, weight modulation [27] in a shared feature transformation for different nodes is introduced:
$\mathbf{H}^{\prime}=\sigma((\mathbf{M}\odot(\mathbf{WH}))\tilde{\mathbf{A}})$ (4)
$\mathbf{h}_{i}^{\prime}=\sigma(\sum_{j\in\mathcal{N}_{i}}{(}\mathbf{m}_{j}\odot\mathbf{W})\mathbf{h}_{j}\tilde{a}_{ij})$ (5)
And then, to enhance feature reuse in GCN and prevent falling into local optima, RSNet [13] implements residual regular connections across different layers:
$\mathbf{H}^{\prime}=\sigma\left(\tilde{\mathbf{A}}((\mathbf{H}\mathbf{W})\odot\mathbf{M})+\mathbf{X}\tilde{\mathbf{W}}\right)$ (6)
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}{\tilde{a}_{ij}}((\mathbf{h}_{j}\mathbf{W})\odot\mathbf{m}_{j})+\mathbf{x}_{i}\widetilde{\mathbf{W}}\right)$ (7)
Finally, it is essential to introduce briefly the transformer, which serves as the backbone of our framework. The transformer architecture is composed of a set of transformer layers, each consisting of two components: a self-attention module and a position-wise feed-forward network. Let $\mathbf{H}\in\mathbb{R}^{T\times N\times F}$ denote the input of the self-attention module where $F$ is the hidden dimension and $\mathbf{h}_{t,i}\in\mathbb{R}^{1\times 1\times F}$ is the hidden representation at joint $i$ in the frame $t$. The input $\mathbf{H}$ is projected by three matrices $\mathbf{W}_{Q}\in\mathbb{R}^{d\times d_{K}}$, $\mathbf{W}_{K}\in\mathbb{R}^{d\times d_{K}}$ and $\mathbf{W}_{V}\in\mathbb{R}^{d\times d_{V}}$ to the corresponding representations $Q$, $K$, $V$. The self-attention is then calculated as:
$\mathbf{Q}$ $=\mathbf{HW}_{Q},\mathbf{K}=\mathbf{HW}_{K},\mathbf{V}=\mathbf{HW}_{V}$ (8)
$\mathbf{A}$ $=\frac{\mathbf{QK}^{\top}}{\sqrt{d_{K}}},\mathrm{Attn}\left(\mathbf{H}\right)=\mathrm{softmax}\left(\mathbf{A}\right)\mathbf{V}$
where $\mathbf{A}$ is a matrix capturing the similarity between queries and keys.
III-B Overall Architecture
Figure 1(a) provides an overview of the proposed framework, which mainly comprises three parts: a joint-based embedding, spatio-temporal criss-cross graph transformer blocks, and a regression head. The joint-based embedding projects the input 2D coordinates of each joint into the feature space. STGFormer blocks aggregate the spatio-temporal context and graph information, updating the representation of each joint. The 3D coordinates are estimated from the learnt features by a regression head.
Figure 1: An overview of our proposed STGFormer. (a) It mainly consists of L sequential STGFormer blocks. (b) The architecture of our spatio-temporal criss-cross graph attention. (c) The architecture of our dual-path modulated hop-wise regular GCN module.
Joint-based embedding. Given a 2D pose sequence represented as $\mathbf{P}_{2D}\in\mathbb{R}^{T\times N\times 2}$, with $T$ and $N$ signifying the number of frames and body joints per frame respectively, we initially map $\mathbf{P}_{2D}$ to high-dimensional embeddings using a joint-based embedding layer. This layer independently processes each 2D coordinate by applying a fully connected layer, followed by a GELU activation. Consequently, the joint-based embedding layer generates features of dimension $T\times N\times F$. It is important to note that in the previous transformer, the embedding layer consolidates all joint coordinates in each frame into a singular vector. This approach reduces the computational demands of subsequent transformer blocks, yet it compromises spatial discrimination. In contrast, our method retains the spatial dimension $N$, while still reducing computational costs through the use of spatio-temporal criss-cross attention.
Spatio-Temporal GraphFormer. The STGFormer block is derived from the traditional transformer block, replacing the original multi-head self-attention layer with spatio-temporal criss-cross graph attention (STG attention). Additionally, a new dual-path modulated hop-wise regular GCN (MHR-GCN) is integrated into the STGFormer block to enhance the descriptive capability of local structures. Detailed discussions on the STG attention and MHR-GCN are provided in Sections III-C and III-D, respectively.
Regression head. A linear regression head is implemented after the STG blocks to predict the 3D pose coordinates $\hat{\mathbf{P}}_{3D}\in\mathbb{R}^{T\times N\times 3}$. The entire architecture is optimized by minimizing the Mean Squared Error (MSE) between the predicted 3D coordinates $\mathbf{\hat{P}}_{3D}$ and the ground-truth 3D coordinates $\mathbf{P}_{3D}$:
$\mathcal{L}=\left\|\hat{\mathbf{P}}_{3D}-\mathbf{P}_{3D}\right\|^{2}$ (9)
III-C Spatio-temporal Criss-cross Graph Attention
Our attention mechanism is specifically designed to effectively model the dependencies between joints. Inspired by [25], we capture spatial and temporal contexts in parallel across different channels, thereby avoiding the quadratic computational cost associated with full attention. Specifically, our mechanism encodes spatio-temporal graph structural relationships into the attention layer to capture structural information between nodes. While modeling the spatial and temporal relationships of the human body, a learnable embedding is allocated based on their structural relationships.
To provide clarity, as shown in figure 1(b), the input embedding $\mathbf{X}\in\mathbb{R}^{T\times N\times F}$ is initially mapped to queries $\mathbf{Q}\in\mathbb{R}^{T\times N\times F}$, keys $\mathbf{K}\in\mathbb{R}^{T\times N\times F}$, and values $\mathbf{V}\in\mathbb{R}^{T\times N\times F}$, all sharing the same dimensions. These are subsequently split into two equal groups along the channel dimension. For clarity in notation, we refer to these split matrices as the time group $\{\mathbf{Q}_{T},\mathbf{K}_{T},\mathbf{V}_{T}\}$ and the space group $\{\mathbf{Q}_{S},\mathbf{K}_{S},\mathbf{V}_{S}\}$. Following this, separate self-attention modules are utilized to calculate the temporal and spatial correlations.
Previous models necessitated the explicit specification of distinct positions or the encoding of positional dependencies within layers. For sequential data, embeddings can be assigned to each position (i.e., absolute position encoding [37]) or encode the relative distance between any two positions in Transformer layers (i.e., relative position encoding [38]). However, for graphs, nodes are not sequentially arranged but are located in multidimensional spaces and connected by edges. To effectively incorporate the structural information of graphs into the model, we introduce a novel spatial encoding technique.
Specifically, we consider a function that measures the spatio-temporal relationships among elements within the graph, which is defined through the connectivity between nodes. We depict the temporal and spatial connectivity relationships of the human body structure in continuous sequences through graphs, and encode these relationships as inputs for spatio-temporal attention on multiple channels in parallel. For temporal encoding and spatial encoding, they are each respectively represented as follows:
$A_{ij}^{t}=\frac{(h_{i}^{t}W_{Q}^{t})(h_{j}^{t}W_{K}^{t})^{T}}{\sqrt{d}}+{b^{t}}_{\psi(v_{i},v_{j})}$ (10)
$A_{ij}^{s}=\frac{(h_{i}^{s}W_{Q}^{s})(h_{j}^{s}W_{K}^{s})^{T}}{\sqrt{d}}+{b^{s}}_{\phi(u_{i},u_{j})}$ (11)
where ${b^{t}}_{\psi(v_{i},v_{j})}$ and ${b^{s}}_{\phi(u_{i},u_{j})}$ are learnable scalars, each indexed by their respective functions $\psi(v_{i},v_{j})$ and $\phi(u_{i},u_{j})$.
Consequently, the above two correlation modules operate in parallel and adhere to the self-attention scheme for feature contextualization. They calculate token-to-token affinities by contextualizing from specific axial perspectives, and effectively complement each other. Finally, we concatenate the outputs from both attention layers along the channel dimension:
$\mathbf{H}_{T}$ $=STGA_{T}(\mathbf{Q}_{T},\mathbf{K}_{T},\mathbf{V}_{T})$ (12)
$\mathbf{H}_{S}$ $=STGA_{S}(\mathbf{Q}_{S},\mathbf{K}_{S},\mathbf{V}_{S})$
$\mathbf{H}$ $=concat\left(\mathbf{H}_{T},\mathbf{H}_{S}\right)$
where $concat$ performs the concatenation. The resultant receptive field of STGA resembles a criss-cross of spatial and temporal axes, and stacking multiple STGA blocks can approximate complete spatio-temporal attention.
III-D Dual-path Modulated Hop-wise Regular GCN Module
To advance the understanding of spatio-temporal relationships among data by directly learning their interconnections, this paper introduces a novel GCN module. Our dual-path modulated hop-wise regular GCN module splits the spatio-temporal graph convolution to independently learn their features, capturing high-order connections and long-range dependencies through an advanced hop-wise approach. Moreover, it employs residual connections to mitigate gradient vanishing and facilitates iterative reuse of features.
As shown in Figure 2, we separate the temporal and spatial graph learning into parallel paths to minimize interference between time and space dimensions. In spatial modulation, a replica of the graph is created for each time step, while in temporal modulation, edges representing temporal relations are introduced to capture dependencies across time, effectively integrating spatial and temporal information.
$\tilde{\mathbf{H}}_{j}^{(\ell)}=$ $\tilde{\mathbf{A}}_{j}(({\mathbf{H}_{S}}^{(\ell)}\mathbf{W}_{j}^{(\ell)})\odot\mathbf{M}_{j}^{(\ell)})$ (13)
$\tilde{\mathbf{H}}_{k}^{(\ell)}=$ $\tilde{\mathbf{A}}_{k}(({\mathbf{H}_{T}}^{(\ell)}\mathbf{W}_{k}^{(\ell)})\odot\mathbf{M}_{k}^{(\ell)})$
where $\mathbf{A}_{j}$ and $\mathbf{A}_{k}$ are the spatial and temporal adjacency matrices, respectively, and $\mathbf{W}_{j}^{(\ell)}$, $\mathbf{W}_{k}^{(\ell)}$ are the corresponding learnable weight matrices. The adjacency modulation matrices $\mathbf{M}_{k}^{(\ell)}$ and $\mathbf{M}_{j}^{(\ell)}$ allow the model presented in this paper to respond flexibly to data variations.
[…]
As illustrated in Figure 3, the concept of hop-wise is introduced in MHR-GCN to capture high-order connection information and long-range dependencies. We use k-hop temporal and j-hop spatial neighbors as shown in figure 1(c). Inspired by the RSNet [13], residual connections are employed in both the temporal and spatial paths to prevent gradient vanishing and bolster feature reuse. The spatio-temporal layer-wise propagation rule of the regular dual-path hop-wise graph network with weight and adjacency modulation is given by
[…]
where $\mathbf{X}_{S}{\tilde{\mathbf{W}}_{j}}^{(\ell)}$ and $\mathbf{X}_{T}{\tilde{\mathbf{W}}_{k}}^{(\ell)}$ are the regular connection, $\sigma\left(\cdot\right)$ is an element-wise nonlinear activation function such as the Gaussian Error Linear Unit (GELU), $\boxplus$ denotes concatenation, ${\tilde{\mathbf{W}}_{j}}^{(\ell)}\in\mathbb{R}^{F\times F}$ and ${\tilde{\mathbf{W}}_{k}}^{(\ell)}\in\mathbb{R}^{F\times F}$ are the learnable weight matrices, $\mathbf{H}_{S}^{(\ell)}\in\mathbb{R}^{N\times F}$ and $\mathbf{H}_{T}^{(\ell)}\in\mathbb{R}^{N\times F}$ are the input feature matrix of the $\ell$-th layer, $\mathbf{H}_{S}^{(\ell+1)}\in\mathbb{R}^{N\times F}$, $\mathbf{H}_{T}^{(\ell+1)}\in\mathbb{R}^{N\times F}$ are the output feature matrix. Thus, the equation of the $\ell+1$ layer feature is reformulated as
[…]
Figure 3: Illustration of the hop-wise concatenation in MHR-GCN for $K=3$ and $J=3$.
IV Experiments
In this section, we first introduce the two benchmark datasets i.e., Human3.6M [39] and MPI-INF-3DHP [40], along with the evaluation metrics used to assess the proposed framework and the experimental setup. Subsequently, we evaluate the proposed framework and compare it with the current state-of-the-art methods, presenting both quantitative and qualitative results. Finally, we verify the impact of each proposed design through ablation studies.
IV-A Experimental Setup
1) Datasets and Evaluation Metrics:
Human3.6M [39] is the most widely used dataset in the evaluation of 3D human pose estimation. It encompasses a vast collection of 3.6 million poses captured using RGB and ToF cameras from diverse viewpoints within a real-world setting. Adhering to the widely used standard protocol, we set five subjects (S1, S5, S6, S7, and S8) for training, and two subjects (S9 and S11) for evaluation. We use Mean Per Joint Position Error (MPJPE) to measure the error under two protocols: Protocol 1 (referred to as P1) and Protocol 2 (referred to as P2). P1 involves calculating MPJPE between the estimated pose and the ground truth after aligning their root joints; P2 involves computing Procrustes-MPJPE (PA-MPJPE), where the ground truth and the pose prediction are further aligned through a rigid transformation.
MPI-INF-3DHP [40] is a large-scale dataset that offers over 2K videos featuring joint annotations of 13 keypoints in outdoor scenes, suitable for 2D and 3D human pose estimation. The ground truth was obtained using a multi-camera arrangement and a marker-less MoCap system. Here, we report the MPJPE, Percentage of Correct Keypoints (PCK) at a threshold of 150mm, and Area Under the Curve (AUC) results as the evaluation metrics.
2) Baseline Methods:
We evaluate the performance of our model by comparing it with state-of-the-art video-based 3D human pose estimation methods, including transformer-based approaches such as StridedFormer [22], MHFormer [21], MixSTE [20], P-STMO [41], HSTFormer [18], Einfalt et al. [19], and PoseFormerV2 [16], as well as GCN-based methods like UGCN [42] and GLA-GCN [11]. Additionally, we also compare our model with methods that combine graphs and transformers, such as STCFormer [25] and MotionAGFormer [43].
3) Implementation Details:
In the experiments, we utilize two kinds of 2D pose sequences as input including the estimated 2D pose by the pre-trained CPN [44] and the real 2D pose (ground truth). For model training, the network parameters are optimized for 40 epochs using the Adam optimizer, with a batch size of 128. The basic learning rate is 0.001, with a decay of 0.97 per epoch. We instantiate our architecture with L = 6 for transformer layers, F = 256 for feature embedding dimensions, H = 8 for self-attention heads, N = 2 for MHR-GCN layers. Our model is implemented using the PyTorch toolkit, version 2.1.0, and runs on RTX 4090 GPU with driver version 525.105.17 and CUDA version 12.0.
IV-B Results and Analysis
1) Quantitative Results:
Performance Comparison on Human3.6M. We conducted a comparative analysis of our method against state-of-the-art approaches on the Human3.6M dataset. Tables I and II present our performance in terms of P1 and P2 errors, respectively, utilizing CPN pretrain 2D coordinates as inputs. Overall, our method has established a new state-of-the-art in a 243-frame setting, outperforming transformer-based, GCN-based, and graph-augmented transformer methods. Benefiting from our strong learning capability for body information, our approach demonstrated improvements over identical transformer-based methods such as StridedFormer [22], MHFormer [21], and MixSTE [20], with gains of 3.4mm, 2.7mm, and 0.6mm on P1 error, and 3.5mm, 2.7mm, and 0.9mm on P2 error, respectively. When compared to UGCN [42], GLA-GCN [11], and STCFormer [25], our method showed enhancements of 5.3mm, 4.1mm, and 0.7mm on P1 error. It is noteworthy that in a series of actions, 10 categories achieved the best-recorded performance to date.
Table III demonstrates the performance of our method on P1 error when using ground-truth 2D poses as input. Compared to CPN 2D poses, ground-truth 2D poses eliminate noise interference, allowing each model to achieve peak performance with ground-truth inputs. Overall, the trend of P1 error changes remains consistent with our method when compared to other methods. Our method reaches state-of-the-art performance at 243 frames input, with a P1 error of 21.4mm, leading the currently most advanced model, MixSTE [20], by 0.2mm.
Performance Comparison on MPI-INF-3DHP. To validate the generalizability of our proposed framework, we tested it on the MPI-INF-3DHP dataset, known for its complex backgrounds. Given the shorter video sequences, we configured the number of input frames at 9, 27, and 81, respectively. Detailed performance comparisons are presented in Table IV. Consistent with the results observed on Human3.6M, the optimal performance is achieved when the frame count is set to 81, achieving PCK of 98.8, AUC of 84.1, and P1 error of 19.2mm. This performance surpasses the current state-of-the-art model by 0.1 in PCK and 0.2 in AUC. These findings highlight the efficacy of the proposed method in handling scenarios of increased complexity.
TABLE I: Performance Comparison on Human3.6M utilizes 2D poses estimated by CPN as inputs under Protocol 1, where Mark * denotes transformer-based methods, Mark ^indicates GCN-based methods, and Mark $\dagger$ represents graph-augmented transformer methods.
Method Publication Dir. Disc. Eat. Greet Phone Photo Pose Purch. Sit SitD. Smoke Wait WalkD. Walk WlkT. Avg.
StridedFormer [22]* (T=243) TMM’22 40.3 43.3 40.2 42.3 45.6 52.3 41.8 40.5 55.9 60.6 44.2 43.0 44.2 30.0 30.2 43.7
MHFormer [21]* (T=351) CVPR’22 39.2 43.1 40.1 40.9 44.9 51.2 40.6 41.3 53.5 60.3 43.7 41.1 43.8 29.8 30.6 43.0
MixSTE [20]* (T=243) CVPR’22 37.6 40.9 37.3 39.7 42.3 49.9 40.1 39.8 51.7 55.0 42.1 39.8 41.0 27.9 27.9 40.9
P-STMO [41]* (T=243) ECCV’22 38.9 42.7 40.4 41.1 45.6 49.7 40.9 39.9 55.5 59.4 44.9 42.2 42.7 29.4 29.4 42.8
HSTFormer [18]* (T=81) arXiv’23 39.5 42.0 39.9 40.8 44.4 50.9 40.9 41.3 54.7 58.8 43.6 40.7 43.4 30.1 30.4 42.7
Einfalt et al. [19]* (T=351) WACV’23 39.6 43.8 40.2 42.4 46.5 53.9 42.3 42.5 55.7 62.3 45.1 43.0 44.7 30.1 30.8 44.2
UGCN [42]^(T=96) ECCV’20 41.3 43.9 44.0 42.2 48.0 57.1 42.2 43.2 57.3 61.3 47.0 43.5 47.0 32.6 31.8 45.6
GLA-GCN [11]^(T=243) CVPR’23 41.3 44.3 40.8 41.8 45.9 54.1 42.1 41.5 57.8 62.9 45.0 42.8 45.9 29.4 29.9 44.4
STCFormer [25$\dagger$ (T=243) CVPR’23 39.6 41.6 37.4 38.8 43.1 51.1 39.1 39.7 51.4 57.4 41.8 38.5 40.7 27.1 28.6 41.0
STGFormer (Ours, T=81) 39.5 41.8 37.2 39.2 43.2 50.5 40.3 40.2 51.1 56.6 42.3 39.0 40.3 27.7 28.1 41.1
STGFormer (Ours, T=243) 38.4 40.7 37.1 38.1 42.7 49.9 39.0 39.3 50.4 56.0 41.4 38.3 39.1 26.9 28.1 40.3
TABLE II: Performance Comparison on Human3.6M utilizes 2D poses estimated by CPN as inputs under Protocol 2, where Mark * denotes transformer-based methods, Mark ^indicates GCN-based methods, and Mark $\dagger$ represents graph-augmented transformer methods.
Method Publication Dir. Disc. Eat. Greet Phone Photo Pose Purch. Sit SitD. Smoke Wait WalkD. Walk WlkT. Avg.
StridedFormer [22]* (T =243) TMM’22 32.7 35.5 32.5 35.4 35.9 41.6 33.0 31.9 45.1 50.1 36.3 33.5 35.1 23.9 25.0 35.2
MHFormer [21]* (T=351) CVPR’22 31.5 34.9 32.8 33.6 35.3 39.6 32.0 32.2 43.5 48.7 36.4 32.6 34.3 23.9 25.1 34.4
MixSTE [20]* (T=243) CVPR’22 30.8 33.1 30.3 31.8 33.1 39.1 31.1 30.5 42.5 44.5 34.0 30.8 32.7 22.1 22.9 32.6
P-STMO [41]* (T=243) ECCV’22 31.3 35.2 32.9 33.9 35.4 39.3 32.5 31.5 44.6 48.2 36.3 32.9 34.4 23.8 23.9 34.4
HSTFormer [18]* (T=81) arXiv’23 31.1 33.7 33.0 33.2 33.6 38.8 31.9 31.5 43.7 46.3 35.7 31.5 33.1 24.2 24.5 33.7
Einfalt et al. [19]* (T=351) WACV’23 32.7 36.1 33.4 36.0 36.1 42.0 33.3 33.1 45.4 50.7 37.0 34.1 35.9 24.4 25.4 35.7
UGCN [42]^(T=96) ECCV’20 32.9 35.2 35.6 34.4 36.4 42.7 31.2 32.5 45.6 50.2 37.3 32.8 36.3 26.0 23.9 35.5
GLA-GCN [11]^(T=243) CVPR’23 32.4 35.3 32.6 34.2 35.0 42.1 32.1 31.9 45.5 49.5 36.1 32.4 35.6 23.5 24.7 34.8
STCFormer [25$\dagger$ (T=243) CVPR’23 29.5 33.2 30.6 31.0 33.0 38.0 30.4 29.4 41.8 45.2 33.6 29.5 31.6 21.3 22.6 32.0
STGFormer (Ours, T=81) 29.6 33.2 30.2 31.3 33.0 37.7 30.1 30.4 41.7 44.6 33.5 29.5 31.7 22.0 23.1 32.1
STGFormer (Ours, T=243) 29.3 33.0 29.9 30.6 32.5 37.5 30.3 29.6 41.0 44.1 33.2 29.1 30.9 21.7 22.8 31.7
TABLE III: Performance Comparison on Human3.6M takes the ground-truth 2D pose as inputs under Protocol 1, where Mark * denotes transformer-based methods, Mark ^indicates GCN-based methods, and Mark $\dagger$ represents graph-augmented transformer methods.
Method Publication Dir. Disc. Eat. Greet Phone Photo Pose Purch. Sit SitD. Smoke Wait WalkD. Walk WlkT. Avg.
StridedFormer [22]* (T =243) TMM’22 27.1 29.4 26.5 27.1 28.6 33.0 30.7 26.8 38.2 34.7 29.1 29.8 26.8 19.1 19.8 28.5
MHFormer [21]* (T=351) CVPR’22 27.7 32.1 29.1 28.9 30.0 33.9 33.0 31.2 37.0 39.3 30.0 31.0 29.4 22.2 23.0 30.5
MixSTE [20]* (T=243) CVPR’22 21.6 22.0 20.4 21.0 20.8 24.3 24.7 21.9 26.9 24.9 21.1 21.5 20.8 14.7 15.7 21.6
P-STMO [41]* (T=243) ECCV’22 28.5 30.1 28.6 27.9 29.8 33.2 31.3 27.8 36.0 37.4 29.7 29.5 28.1 21.0 21.0 29.3
HSTFormer [18]* (T=81) arXiv’23 24.9 27.4 28.1 25.9 28.2 33.5 28.9 26.8 33.4 38.2 27.2 26.7 27.1 20.4 20.8 27.8
GLA-GCN [11]^(T=243) CVPR’23 26.5 27.2 29.2 25.4 28.2 31.7 29.5 26.9 37.8 39.9 29.9 27.0 27.3 20.5 20.8 28.5
STCFormer [25$\dagger$ (T=243) CVPR’23 21.4 22.6 21.0 21.3 23.8 26.0 24.2 20.0 28.9 28.0 22.3 21.4 20.1 14.2 15.0 22.0
STGFormer (Ours, T=81) 22.6 22.3 21.5 21.2 23.6 25.3 24.8 19.9 27.9 28.0 22.9 21.5 20.4 14.5 16.8 22.2
STGFormer (Ours, T=243) 21.2 21.7 20.6 20.6 22.7 24.0 24.4 19.6 26.4 27.1 22.6 21.3 19.6 14.1 15.3 21.4
TABLE IV: Performance Comparison on MPI-INF-3DHP: Mark * denotes transformer-based methods, Mark ^indicates GCN-based methods, and Mark $\dagger$ represents graph-augmented transformer methods.
[…]
MHFormer [21]* (T=9) CVPR’22 93.8 63.3 58.0
MixSTE [20]* (T=27) CVPR’22 94.4 66.5 54.9
P-STMO [41]* (T=81) ECCV’22 97.9 75.8 32.2
HSTFormer [18]* (T=81) arXiv’23 98.0 78.6 28.3
PoseFormerV2 [16]* (T =81) CVPR’23 97.9 78.8 27.8
Einfalt et al. [19]* (T=81) WACV’23 95.4 67.6 46.9
UGCN [42]^(T=96) ECCV’20 86.9 62.1 68.1
GLA-GCN [11]^(T=81) CVPR’23 98.5 79.1 27.8
STCFormer [25$\dagger$ (T=81) CVPR’23 98.7 83.9 23.1
MotionAGFormer [43$\dagger$ (T=27) WACV’23 98.2 83.5 19.2
STGFormer (Ours, T=27) 98.2 83.8 22.6
STGFormer (Ours, T=81) 98.8 84.1 19.2
2) Qualitative Results:
In this section, we conduct a qualitative analysis of the model’s performance by visualizing the attention layers and assessing the effects of model on Human3.6M dataset.
Attention visualization. We visualized the final spatial and temporal attention layers in our module, as shown in Figure 3. As expected, Figure 4(a) shows that the spatial attention layer map indicates that our proposed model has learned various patterns of different joints during movements. On the temporal attention layer, Figure 4(b) reveals a strong correlation between joints across adjacent frames and higher-order frames.
Figure 4: Visualization of attention layer maps from the spatial and temporal attention modules in the STGFormer.
Result visualization on Human3.6M. The visualization of our model’s performance on the Human3.6M dataset is depicted in Figure 5, where four examples from the Walking, Sitting, Greeting, and Eating sequences were randomly selected. Due to the framework’s strong modeling capabilities in both temporal and spatial dimensions for human action sequences, our method exhibits robust performance in complex actions such as Sitting (second row) and Eating (fourth row), effectively reconstructing poses in three-dimensional space.
Figure 5: Examples of 3D pose estimation by our framework on Human3.6M dataset.
3) Ablation Studies
To further investigate the proposed model, we conducted a series of ablation studies on the Human3.6M dataset using CPN-pretrained 2D keypoint coordinates as input. The first set of ablation studies focused on how the performance of our proposed model varies under different frame counts. As illustrated in Table V, the P1 error comparison of our model with other methods on the Human3.6M dataset is detailed for 27, 81, and 243 frames. Overall, we observed a trend of decreasing P1 error with an increasing frame count in our method and others. Among these methods, our model achieved the best results at 27, 81, and 243 frames, demonstrating strong performance in 3D pose estimation in long sequence videos. Notably, with a comparable number of parameters, our model requires less than 79% of the FLOPs compared to the currently best-performing MixSTE [20] at 243 frames.
TABLE V: The P1 error comparisons with different numbers of frame on Human3.6M dataset. The best result in each column is highlighted in red.
Method Frames Parameters FLOPs (M) P1 (mm)
StridedFormer [22] 27 4.01M 163 46.9
MHFormer [21] 27 18.92M 1000 45.9
MixSTE [20] 27 33.61M 15402 45.1
P-STMO [41] 27 4.6M 164 46.1
STCFormer [25] 27 4.75M 2173 44.1
STGFormer 27 12.66M 3104 43.7
StridedFormer [22] 81 4.06M 392 45.4
MHFormer [21] 81 19.67M 1561 44.5
MixSTE [20] 81 33.61M 46208 42.7
P-STMO [41] 81 5.4M 493 44.1
STCFormer [25] 81 4.75M 6520 42.0
STGFormer 81 20.08M 9632 41.1
StridedFormer [22] 243 4.23M 1372 44.0
MHFormer [21] 243 24.72M 4812 43.2
MixSTE [20] 243 33.61M 138623 40.9
P-STMO [41] 243 6.7M 1737 42.8
STCFormer [25] 243 4.75M 19561 41.0
STGFormer 243 31.91M 28871 40.3
The second set of ablation studies examined the effects of various components within our proposed model on its performance, utilizing CPN-pretrained 2D coordinate data from 27 frames as input. Initially, spatial graph attention (SGA) and temporal graph attention (TGA) were analyzed separately, with MHR-GCN components replaced by an MLP. Subsequently, only the combined spatial-temporal graph attention was operational in the STGA-only set. Spatial MHR-GCN (SMHR-GCN) and Temporal MHR-GCN (TMHR-GCN) constitute the MHR-GCN components. Table VI offers detailed insights into how each component influences the overall performance of our model. Employing either spatial or temporal graph attention alone resulted in a substantial increase in P1 error, reaching 263.2mm and 57.8mm, respectively. By integrating both spatial and temporal graph attention, and adding SMHR-GCN and TMHR-GCN, the errors were further reduced by 6.4mm and 8.3mm, respectively.
TABLE VI: Performance contribution of each component in the proposed framework on Human3.6M dataset.
STGA SMHR-GCN TMHR-GCN P1 (mm)
SGA 263.2
TGA 57.8
STGA $\checkmark$ 53.3
+SMHR-GCN $\checkmark$ $\checkmark$ 46.9
+TMHR-GCN $\checkmark$ $\checkmark$ 45.0
+MHR-GCN $\checkmark$ $\checkmark$ $\checkmark$ 43.7
V Conclusion
In this paper, we introduce the Spatio-Temporal GraphFormer, a novel model that integrates long-range temporal-spatial dependencies with graph structural information in its attention mechanisms and utilizes a dual-path modulated hop-wise regular GCN Module to enhance the extraction of graph information for 3D human pose estimation in videos. Specifically, the spatio-temporal criss-cross graph attention separately models spatial and temporal correlations, leveraging encoded graph structure information within the model. The dual-path modulated hop-wise regular GCN Module adjusts graph weights and accounts for high-order dependencies across both time and space dimensions. Experiments on two benchmarks affirm the efficacy and superior generalization capability of our proposed model over state-of-the-art techniques. For future work, we plan to develop a multi-person skeletal sequence spatio-temporal graph representation method to extend our model’s applicability to 3D human pose estimation in multi-person scenarios and broaden its use in various contexts.